An electron is projected into an electric field An electron is projected into an electric field &n at an angle of 125° with respect to the x-axis with speed 1525 m/s. What is the acceleration of the electron once it is in the electric field (neglect any gravitational effects)? You may consider the electron to enter the field at the origin of coordinates. What is the position of the electron after 15 s, assuming it remains in the same field for the whole time?

Full solution please!
Answer #1


electric force, Fe = qE

and a = Fe/m = qE/m

a = (1.6 x 10^-19 x 3.2 ) / (9.109 x 10^-31)

a = 5.62 x 10^11 m/s^2 (along -ve x axis)   .....Ans


Applying r = ut + at^2/2

u = 1520 (cos125i + sin125j) = -871.84i + 1245.11j m/s


r = (-871.84i + 1245.11j)(15) + (15^2/2) (- 5.62 x 10^11 i )

r = -13077.6i + 18676.65j - 6.32x10^13i

r = - 6.32 x 10^13i + 18676.65j m

so electrons corrdinates are ( -6.32 x 10^13, 18676.65) m